OpenStudy (anonymous):

simplify \[\sqrt{3}(\sqrt{12}-\sqrt{6})\]

OpenStudy (lalaly):

\[\sqrt{12}= 2\sqrt3\] \[\sqrt{3}(2 \sqrt3-\sqrt6)\]\[=6-\sqrt{18}\]\[=6-3\sqrt2\]